A says to B " I am twice as old as you were when I was as old as you are".The sum of their ages be 63 yrs then find their present ages?
Answer by J2R2R(94)

A says to B “I am twice as old as you were when I was as old as you are". The sum of their ages is 63 yrs then find their present ages?

When A was what B’s age is now, let X be the age of B, Y be the gap between the ages of A and B, and so A’s age was X + Y.

They have now moved on Y years so that B is now what A was.

Now B’s age is X + Y and A’s age is X + 2Y, but this also equals 2X by the statement “I am twice as old as you were when I was as old as you are”

So, X + 2Y = 2X, giving X = 2Y

Therefore A’s age was X + Y = 2Y + Y = 3Y and B’s age was X = 2Y

Now A’s age is 3Y + Y = 4Y and B’s age is 2Y + Y = 3Y

The present ages of A and B total 63, so 4Y + 3Y = 7Y = 63 which gives Y = 9 (the gap)

Since X = 2Y = 18, we have the solution.

A was 3Y = 27 and B was 2Y = 18 (9 years gap)

A is now 4Y = 36 and B is now 3Y = 27 (9 years gap)

A is now (36) twice what B was (18) when A was what B is now (27).